# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ 4 Numerical Experiments
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE CP AIL
@@
$n=400,r=n^{0.8}$0.27 0.38(0.35) -0.32 0.43(0.44) 0.52 0.43 92.9% 1.86(1.14)
$r=n^{0.85}$0.20 0.34(0.33) -0.24 0.40(0.41) 0.48 0.43 94.8% 2.08(1.31)
$r=n^{0.9}$0.13 0.32(0.31) -0.16 0.39(0.40) 0.45 0.45 96.2% 2.30(1.45)
Boostrap with $f_{0}$0.08 0.46(0.41) -0.01 0.63(0.65) 0.62 - 92.4% 4.46(3.77)
$n=700,r=n^{0.8}$0.10 0.29(0.30) -0.20 0.37(0.38) 0.42 0.38 94.3% 1.79(0.95)
$r=n^{0.85}$0.06 0.28(0.29) -0.15 0.35(0.36) 0.40 0.38 95.8% 1.92(1.02)
$r=n^{0.9}$ 0.02 0.27(0.29) -0.06 0.33(0.34) 0.40 0.39 97.2% 2.18(1.22)
Boostrap with $f_{0}$0.02 0.41(0.38) -0.11 0.55(0.56) 0.56 - 95.0% 4.16(3.10)
Binomial Model
$n=400,r=n^{0.8}$ 0.42 0.65(0.57) 0.04 0.08(0.09) 0.86 0.64 87.6% 0.30(0.00)
$r=n^{0.85}$ 0.32 0.58(0.51) 0.03 0.07(0.08) 0.77 0.66 91.3% 0.31(0.00)
$r=n^{0.9}$ 0.24 0.53(0.45) 0.03 0.07(0.07) 0.70 0.68 94.7% 0.33(0.00)
Boostrap with $f_{0}$ 0.25 0.72(0.61) 0.02 0.09(0.10) 0.94 - 95.5% 0.48(0.24)
$n=700,r=n^{0.8}$ 0.24 0.45(0.39) 0.03 0.06(0.06) 0.59 0.55 94.3% 0.27(0.01)
$r=n^{0.85}$ 0.15 0.41(0.35) 0.02 0.05(0.05) 0.54 0.56 96.0% 0.28(0.01)
$r=n^{0.9}$ 0.09 0.39(0.33) 0.02 0.05(0.05) 0.51 0.57 97.3% 0.30(0.01)
Boostrap with $f_{0}$ 0.09 0.60(0.46) 0.00 0.08(0.08) 0.76 - 94.9% 0.39(0.20)
Figure 1: Histograms of $\psi^{\prime}(\widehat{f}^{B}(x_{\text{test}}))$ and $\psi^{\prime}(f_{0}(x_{\text{test}}))$. Results are shown for (1(a)) logistic model, (1(b)) Poisson model and (1(c)) Binomial model with $n=700$ and subsampling ratio $r=n^{0.9}$.
For Aim (i), we randomly select sub-datasets of size $r$ to train neural networks, varying $r$ from $n^{0.8}$ to $n^{0.9}$, with $n=400$ and $700$. Based on our experiments, we recommend $r=n^{0.9}$, balancing two key factors: (1) a sufficiently large $r$ ensures neural network convergence and strong performance by providing ample training data, and (2) $r$ remains smaller than the total sample size $n$. To maintain computational feasibility, we set the number of subsampling iterations to $B=1400$. The testing data, generated similarly to the training data, has a size of $n_{\text{test}}=1000$. We apply the DNN-based ESM algorithm with a network structure of $L=3$ and $\mathbf{p}=(p,128,64,1)$. Each hidden layer incorporates a 10% dropout rate for regularization, and we apply $\ell_{2}$ regularization with a weight decay rate of 0.02. The learning rate is fine-tuned for optimal validation convergence. To keep network outputs within a reasonable range, we implement early stopping with a maximum of 500 epochs to prevent overfitting. The activation function used is ReLU.
With $\widehat{f}^{B}(\cdot)$ being the output of our algorithm, we evaluate its performance using $n_{\text{test}}$ test points across the following metrics.
  • $\text{Bias}_{f}$: the average bias between $\widehat{f}^{B}$ and $f_{0}$.
  • $\text{MAE}_{f}$: the mean absolute error between $\widehat{f}^{B}$ and $f_{0}$.
  • $\text{Bias}_{\psi^{\prime}}$: the average bias in the transformed form $\psi^{\prime}(\widehat{f}^{B})$ and $\psi^{\prime}(f_{0})$.
  • $\text{MAE}_{\psi^{\prime}}$: the mean absolute error between $\psi^{\prime}(\widehat{f}^{B})$ and $\psi^{\prime}(f_{0})$.
  • EmpSD: the empirical standard deviation of $\widehat{f}^{B}$, given by
    $\bigg{(}\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}(\widehat{f}^{B}(\mathbf{x}_{\text{test},i})-f_{0}(\mathbf{x}_{\text{test},i}))^{2}\bigg{)}^{\frac{1}{2}}.$
  • SE: the square root of the average estimated variance $\widehat{\sigma}_{*}^{2}$.
  • CP (Coverage Probability): the proportion of times the inference interval covers the true value $\psi^{\prime}(f_{0}(\mathbf{x}_{\text{test},i}))$ for $i\in[n_{\text{test}}]$.
  • AIL (Average Interval Length): $\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}(U_{i}-L_{i}),$ where $U_{i}$ and $L_{i}$ are the upper and lower bounds of the prediction interval for test instance $\mathbf{x}_{\text{test},i}$.
Numbers in brackets reported in the tables represent the standard deviation of each metric across test data points.
Table 1 presents the results of our method across logistic, Poisson, and Binomial models. Increasing the sample size $n$ reduces prediction deviations, as reflected in decreasing mean absolute errors (MAE) for both $f$ and its transformed counterpart $\psi^{\prime}$. The close alignment between empirical standard deviations (EmpSD) and estimated standard errors (SE) confirms the accuracy of SE. Our method consistently achieves coverage probabilities (CP) near the 95% target. Additionally, the average interval lengths (AIL) are substantially shorter than those from bootstrap intervals, suggesting its efficiency in uncertainty estimation. In the bootstrap approach with unknown $f_{0}$, we estimate $\widehat{f}$ from one data half and evaluate the distribution of $f_{0}-\widehat{f}$ on the other half to determine the 95 percentile. While bootstrap with $f_{0}$ achieves nominal coverage (95%), its wider intervals indicate conservative uncertainty estimates. Figure 1 supports our findings and show that $\psi^{\prime}(\widehat{f}^{B}(\mathbf{x}_{\operatorname{test}}))$ aligns well with $\psi^{\prime}(f_{0}(\mathbf{x}_{\operatorname{test}}))$. Despite the wide range of true values, our confidence intervals remain tight, with AILs as low as 0.43–0.45 for logistic models and 1.86–2.3 for Poisson models, indicating the precision of ESM in uncertainty quantification.
Table 2: Simulation results for different models under varying sample sizes $n$ and subsampling ratios $r$ with non-linear function $g_{1}$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE CP AIL

### v3

@@ 6 Discussion
Our theoretical analysis focuses on ReLU activations, where monomial-based approximations are tractable. We envision that ESM can be extended to networks with smooth activation functions such as tanh or GELU, although this requires careful consideration. Unlike ReLU, smooth activations are less flexible in local adaptations, which may result in slower approximation rates and influence convergence behavior. To preserve inference validity in this setting, it may be necessary to employ alternative analytical tools or apply stronger regularization to control model complexity and prevent saturation. Moreover, our analysis assumes that DNN estimators, trained via gradient descent, effectively minimize empirical risk and generalize well, an increasingly supported view in recent theory. While we do not analyze training dynamics explicitly, our assumptions are consistent with emerging work on implicit bias and convergence in overparameterized models (Chizat & Bach, 2020; Chatterji et al., 2021; Cao et al., 2022). An important direction is to examine how implicit regularization techniques, such as early stopping, dropout, and weight decay, enhance ESM’s applicability by promoting model stability. These methods curb overfitting, reduce sensitivity to data perturbations, and help satisfy the smoothness and exchangeability conditions required for valid U-statistic–based inference.
Data Availability Statement. The data used in this study are publicly available from the eICU Collaborative Research Database (https://eicu-crd.mit.edu/) upon completion of a data use agreement and approval through the PhysioNet credentialed access process.
Supplementary Material for “Inference for Deep Neural Network Estimators in Generalized Nonparametric Models”
Appendix A Additional Numerical Results
We conduct additional numerical experiments to further evaluate our proposed method. In Section A.1, we extend the approach to the binomial regression setting. In Section A.2, we compare neural networks and random forests, showing that their relative performance varies across different scenarios. In Section A.3, we further examine simpler kernel regression models to assess ESM under smoother nonlinear structures. In Section A.4, we present experiments illustrating how our method performs as data dimensionality and nonlinearity increase. We also investigate the effects of key optimization and design parameters, including the subsample size $r$, the number of subsamples $B$, the network depth $L$, and the dropout rate, in Sections A.5, A.6, A.7, A.8, and A.9, respectively.
A.1 Binomial Regression
We show the results under binomial regression, that is, with $k\in\{0,1,...,n_{\text{trial}}\}$,
$\mathbb{P}(y_{i}=k|\mathbf{x}_{i})={\binom{n_{\text{trial}}}{k}}p^{k}\big(\mathbf{x}^{(i)}\big)\big(1-p\big(\mathbf{x}^{(i)}\big)\big)^{k},$
with $p\big(\mathbf{x}^{(i)}\big)=\frac{1}{1+\exp\{-g(\mathbf{x}_{i})\}}.$ Under this model, $f_{0}(\mathbf{x})=g(\mathbf{x})$ and $\psi^{\prime}(x)=n_{\text{trial}}(1+\exp(-x))^{-1}$.
We use the same simulation framework as described in Section 4. We set the number of trials in the binomial distribution to $n_{\text{trial}}=5$, while keeping all other simulation parameters and settings consistent with the baseline experiments. At the same time, we also vary the resampling number $B$ from 1400 to 3000, and present the corresponding results to examine the effect of $B$ on the variance estimation and interval construction.
Table S.1: Simulation results for Binomial regression under varying sample sizes $n$ and subsampling ratios $r$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
$B=1400$
$n=400,r=n^{0.8}$ 0.03(0.34) 0.27(0.21) 0.01(0.07) 0.06(0.04) 0.31 0.39 0.32 94.0% 0.26(0.07)
$r=n^{0.85}$ 0.04(0.36) 0.28(0.23) 0.01(0.07) 0.06(0.05) 0.34 0.41 0.34 93.5% 0.27(0.08)
$r=n^{0.9}$ 0.04(0.34) 0.27(0.21) 0.01(0.07) 0.06(0.04) 0.32 0.43 0.35 95.0% 0.28(0.08)
$n=700,r=n^{0.8}$ 0.02(0.27) 0.21(0.17) 0.00(0.06) 0.05(0.03) 0.25 0.34 0.25 93.2% 0.21(0.05)
$r=n^{0.85}$ 0.03(0.25) 0.20(0.16) 0.01(0.05) 0.04(0.03) 0.24 0.34 0.25 94.3% 0.21(0.05)
$r=n^{0.9}$ 0.03(0.26) 0.20(0.16) 0.01(0.06) 0.04(0.03) 0.25 0.34 0.25 93.5% 0.21(0.05)
$B=3000$
$n=400,r=n^{0.8}$ 0.05(0.34) 0.27(0.21) 0.01(0.07) 0.06(0.04) 0.31 0.36 0.32 93.9% 0.26(0.07)
$r=n^{0.85}$ 0.03(0.35) 0.27(0.22) 0.01(0.07) 0.06(0.05) 0.33 0.38 0.34 94.3% 0.27(0.07)
$r=n^{0.9}$ 0.03(0.34) 0.27(0.21) 0.00(0.07) 0.06(0.04) 0.32 0.38 0.34 95.0% 0.28(0.07)
$n=700,r=n^{0.8}$ 0.04(0.26) 0.20(0.16) 0.01(0.06) 0.04(0.03) 0.24 0.29 0.25 93.3% 0.20(0.05)
$r=n^{0.85}$ 0.01(0.25) 0.19(0.15) 0.00(0.05) 0.04(0.03) 0.23 0.29 0.24 94.4% 0.20(0.05)
$r=n^{0.9}$ 0.02(0.24) 0.19(0.15) 0.00(0.05) 0.04(0.03) 0.23 0.29 0.24 94.1% 0.20(0.05)
Figure S.1: Estimation and inference in simulation samples: Binomial Model with $n=700$ and $r=n^{0.9}$. Figure 1(a) and 1(b) show corrected and uncorrected estimation SE versus empirical SD of all test samples with $B=1400$ and $B=3000$, respectively.
As shown in Table S.1, our method achieves reasonable coverage probabilities and stable interval lengths under the binomial regression setting, across various values of $B$ and $r$. In binomial regression, the corrected standard errors $\text{SE}_{c}$ closely track the empirical variability, supporting the validity of our variance estimation procedure. We also plot the corrected and uncorrected estimation SE versus empirical SD with different $B$ in Figure S.1. Figure S.1 also shows that with a larger $B$, the uncorrected estimator becomes more accurate and gradually approaches the empirical SD. In contrast, the corrected SE consistently aligns well with the empirical SD across different values of $B$, demonstrating the effectiveness of our correction procedure, especially in finite-sample settings.
A.2 Comparison to Random Forests
Although structurally different from deep neural networks (DNNs), random forests (RFs) are another widely used tool for nonparametric estimation. As a benchmark, we compare the performance of neural networks and random forests within our simulation framework to examine their respective behaviors across varying settings, rather than to identify a universally superior method. In addition, we investigate the use of our corrected variance estimator within the RF framework to enhance the accuracy of empirical standard deviation estimation.
Table S.2: Comparison under the basic setting in logistic model.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic Model with DNNs
$n=400,r=n^{0.8}$ 0.11(0.72) 0.57(0.44) 0.02(0.14) 0.11(0.08) 0.65 0.78 0.66 91.7% 0.46(0.13)
$r=n^{0.85}$ 0.06(0.59) 0.47(0.36) 0.01(0.12) 0.10(0.07) 0.55 0.68 0.58 94.0% 0.43(0.11)
$r=n^{0.9}$ 0.07(0.50) 0.40(0.31) 0.01(0.10) 0.08(0.06) 0.48 0.59 0.50 94.6% 0.39(0.09)
$n=700,r=n^{0.8}$ 0.08(0.41) 0.33(0.26) 0.02(0.09) 0.07(0.05) 0.39 0.51 0.39 93.6% 0.31(0.08)
$r=n^{0.85}$ 0.05(0.36) 0.28(0.22) 0.01(0.08) 0.06(0.05) 0.34 0.44 0.34 93.9% 0.28(0.06)
$r=n^{0.9}$ 0.06(0.33) 0.26(0.21) 0.01(0.07) 0.06(0.05) 0.31 0.40 0.31 93.5% 0.26(0.06)
Logistic Model with RFs
$n=400,r=n^{0.8}$ 0.07(0.39) 0.31(0.24) 0.02(0.09) 0.07(0.05) 0.31 0.38 0.33 89.1% 0.28(0.03)
$r=n^{0.85}$ 0.07(0.41) 0.33(0.25) 0.02(0.09) 0.07(0.06) 0.36 0.43 0.36 91.1% 0.31(0.04)
$r=n^{0.9}$ 0.08(0.44) 0.35(0.27) 0.02(0.10) 0.08(0.06) 0.40 0.48 0.41 93.0% 0.35(0.05)
$n=700,r=n^{0.8}$ 0.07(0.33) 0.27(0.21) 0.02(0.07) 0.06(0.05) 0.28 0.36 0.28 89.1% 0.25(0.03)
$r=n^{0.85}$ 0.06(0.36) 0.29(0.22) 0.01(0.08) 0.06(0.05) 0.32 0.41 0.32 91.8% 0.27(0.03)
$r=n^{0.9}$ 0.06(0.37) 0.30(0.23) 0.01(0.08) 0.07(0.05) 0.35 0.47 0.37 94.2% 0.31(0.04)
Table S.3: Comparison under the basic setting in Poisson model.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Poisson Model with DNNs
$n=400,r=n^{0.8}$ -0.32(0.38) 0.38(0.31) -0.16(0.23) 0.22(0.18) 0.35 0.43 0.36 88.6% 0.86(0.45)
$r=n^{0.85}$ -0.23(0.36) 0.33(0.27) -0.12(0.25) 0.21(0.18) 0.34 0.43 0.36 91.9% 0.95(0.52)
$r=n^{0.9}$ -0.14(0.34) 0.29(0.24) -0.06(0.25) 0.19(0.18) 0.33 0.42 0.35 94.3% 1.02(0.59)
$n=700,r=n^{0.8}$ -0.17(0.26) 0.25(0.20) -0.10(0.19) 0.16(0.14) 0.25 0.34 0.26 90.7% 0.67(0.34)
$r=n^{0.85}$ -0.10(0.26) 0.21(0.17) -0.06(0.19) 0.15(0.14) 0.24 0.33 0.25 93.2% 0.71(0.38)
$r=n^{0.9}$ -0.04(0.24) 0.19(0.15) -0.02(0.19) 0.14(0.13) 0.23 0.32 0.24 94.5% 0.72(0.38)
Poisson Model with RFs
$n=400,r=n^{0.8}$ -0.03(0.24) 0.19(0.16) -0.01(0.17) 0.13(0.11) 0.21 0.25 0.21 92.4% 0.60(0.21)
$r=n^{0.85}$ -0.03(0.26) 0.20(0.17) 0.00(0.18) 0.14(0.11) 0.24 0.29 0.24 93.7% 0.68(0.25)
$r=n^{0.9}$ -0.02(0.27) 0.21(0.18) 0.01(0.20) 0.15(0.13) 0.26 0.32 0.28 94.6% 0.79(0.30)
$n=700,r=n^{0.8}$ -0.03(0.21) 0.16(0.14) -0.01(0.15) 0.11(0.09) 0.18 0.24 0.19 92.8% 0.52(0.17)
$r=n^{0.85}$ -0.03(0.22) 0.17(0.14) 0.00(0.16) 0.12(0.10) 0.20 0.28 0.21 94.2% 0.59(0.20)
$r=n^{0.9}$ -0.03(0.26) 0.20(0.17) 0.00(0.18) 0.14(0.12) 0.24 0.32 0.25 94.3% 0.69(0.26)
We first present random forest performance under the basic settings as in the simulation section of the main text, covering logistic and Poisson scenarios and using the same evaluation metrics as for neural networks.
Table S.4: Comparison under the changing setting in logistic model.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic Model with DNNs
$n=400,r=n^{0.8}$ 0.07(0.45) 0.36(0.28) 0.01(0.09) 0.07(0.06) 0.42 0.50 0.43 93.8% 0.32(0.09)
$r=n^{0.85}$ 0.04(0.40) 0.32(0.25) 0.01(0.08) 0.07(0.05) 0.38 0.45 0.38 93.9% 0.30(0.08)
$r=n^{0.9}$ 0.04(0.37) 0.29(0.23) 0.01(0.08) 0.06(0.05) 0.34 0.41 0.35 93.8% 0.28(0.07)
$n=700,r=n^{0.8}$ 0.05(0.29) 0.23(0.19) 0.01(0.06) 0.05(0.04) 0.26 0.34 0.27 93.3% 0.22(0.06)
$r=n^{0.85}$ 0.04(0.29) 0.23(0.19) 0.01(0.06) 0.05(0.04) 0.25 0.32 0.25 90.4% 0.21(0.05)
$r=n^{0.9}$ 0.04(0.29) 0.22(0.19) 0.01(0.06) 0.05(0.04) 0.22 0.31 0.23 90.0% 0.19(0.05)
Logistic Model with RFs
$n=400,r=n^{0.8}$ 0.02(0.46) 0.36(0.30) 0.01(0.09) 0.07(0.06) 0.32 0.38 0.32 84.9% 0.28(0.04)
$r=n^{0.85}$ 0.04(0.46) 0.36(0.30) 0.01(0.10) 0.08(0.06) 0.35 0.42 0.36 87.6% 0.31(0.05)
$r=n^{0.9}$ 0.03(0.48) 0.37(0.30) 0.01(0.10) 0.08(0.06) 0.38 0.48 0.41 90.8% 0.34(0.05)
$n=700,r=n^{0.8}$ 0.03(0.40) 0.31(0.26) 0.01(0.08) 0.06(0.05) 0.28 0.36 0.28 84.8% 0.24(0.03)
$r=n^{0.85}$ 0.02(0.41) 0.32(0.26) 0.01(0.08) 0.07(0.05) 0.32 0.41 0.32 88.5% 0.27(0.04)
$r=n^{0.9}$ 0.01(0.42) 0.33(0.27) 0.01(0.09) 0.07(0.05) 0.34 0.47 0.37 91.3% 0.31(0.05)
Table S.5: Comparison under the changing setting in Poisson model.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Poisson Model with DNNs
$n=400,r=n^{0.8}$ -0.20(0.31) 0.28(0.24) -0.12(0.24) 0.19(0.19) 0.29 0.35 0.30 90.8% 0.80(0.52)
$r=n^{0.85}$ -0.12(0.29) 0.24(0.20) -0.07(0.24) 0.17(0.18) 0.27 0.34 0.29 93.3% 0.82(0.58)
$r=n^{0.9}$ -0.06(0.27) 0.21(0.18) -0.03(0.24) 0.16(0.18) 0.25 0.32 0.27 95.0% 0.81(0.59)
$n=700,r=n^{0.8}$ -0.10(0.22) 0.18(0.15) -0.07(0.18) 0.13(0.14) 0.20 0.27 0.20 92.3% 0.59(0.37)
$r=n^{0.85}$ -0.04(0.20) 0.16(0.13) -0.03(0.17) 0.12(0.13) 0.19 0.26 0.19 94.1% 0.59(0.38)
$r=n^{0.9}$ -0.01(0.20) 0.15(0.12) -0.01(0.17) 0.12(0.12) 0.18 0.24 0.18 93.4% 0.58(0.37)
Poisson Model with RFs
$n=400,r=n^{0.8}$ -0.00(0.24) 0.19(0.15) -0.00(0.20) 0.14(0.14) 0.21 0.25 0.21 91.2% 0.62(0.25)
$r=n^{0.85}$ -0.00(0.26) 0.20(0.16) 0.00(0.20) 0.15(0.14) 0.23 0.28 0.24 93.3% 0.71(0.30)
$r=n^{0.9}$ -0.00(0.28) 0.22(0.18) 0.02(0.23) 0.17(0.16) 0.26 0.33 0.28 93.4% 0.84(0.40)
$n=700,r=n^{0.8}$ 0.00(0.21) 0.16(0.14) -0.00(0.17) 0.12(0.12) 0.18 0.24 0.18 91.4% 0.54(0.21)
$r=n^{0.85}$ 0.01(0.22) 0.17(0.14) 0.01(0.18) 0.13(0.12) 0.20 0.27 0.21 93.2% 0.63(0.26)
$r=n^{0.9}$ -0.00(0.26) 0.20(0.16) 0.01(0.20) 0.15(0.13) 0.24 0.32 0.25 94.1% 0.73(0.33)
As shown in Tables S.2 and S.3, when the sample size is small, neural networks tend to produce wider estimation intervals compared to random forests, reflecting greater estimation uncertainty. However, as both the sample size $n$ and subsample size $r$ increase, neural networks are able to capture the signal more effectively, leading to interval lengths and coverage that are comparable to, or even better than those of random forests. The findings are consistent with the modern generalization theory of neural networks (Cao et al. 2022, Kou et al. 2023), which suggests that neural networks can generalize well when data quality is high and sufficient samples are available. In contrast, random forests tend to offer stronger robustness in low-data or high-noise scenarios.
To further elaborate on these points, we conduct additional experiments under a modified simulation setting, where we increase the signal strength and reduce the data noise. Specifically, we generate the covariates $\mathbf{x}_{i}$ from a multivariate normal distribution $\mathcal{N}(0,\frac{1}{4}\mathbf{I}_{10})$, and define the underlying regression function as $g(\mathbf{x})=2x_{1}+x_{2}^{2}+0.1\cdot\arctan(x_{3}-0.3).$ The other settings remain unchanged. This setup introduces weaker noise component and stronger signal component.
The simulation results are summarized in Tables S.4 and S.5. In the logistic regression setting, neural networks tend to yield smaller estimation bias and narrower confidence intervals, while also achieving slightly higher coverage probabilities compared to random forests. This shows that, in scenarios with relatively strong signal structures and adequate subsample sizes, neural networks will offer improved estimation efficiency. For the Poisson model, both methods perform comparably in terms of coverage. However, as the subsample size $r$ increases, the interval lengths produced by neural networks generally decrease, whereas those from random forests tend to increase. Overall, both methods have their strengths: RFs offer stability and robustness under noise, while DNNs can capture more complex patterns when data quality allows.
Finally, in all of these settings, we find that the corrected variance formula performs well for random forests. The corrected standard errors align more closely with the empirical standard deviations compared to the uncorrected ones. As a result, the coverage probabilities achieved by random forests are close to the nominal level.
A.3 Additional Experiments on Kernel Regressions
We consider a kernelized generalized linear model (GLM) to evaluate ESM under smooth nonlinear structures. Let $\gamma>0$ and $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\{-\gamma\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}\}$ denote the Gaussian kernel. To make the kernel GLM computationally tractable for moderate to large sample sizes, we adopt the Nystrom approximation to obtain a finite dimensional feature representation $\phi(\mathbf{x})\in\mathbb{R}^{m}$, where $k(\mathbf{x},\mathbf{x}^{\prime})$ can be approximated by $k(\mathbf{x},\mathbf{x}^{\prime})\approx\phi(\mathbf{x})^{\top}\phi(\mathbf{x}^{\prime})$. Under this approximation, the kernel function becomes $f(\mathbf{x})\approx\mathbf{w}^{\top}\phi(\mathbf{x})$, and the regularized empirical loss reduces to
$\widehat{\mathbf{w}}=\operatorname{argmin}_{\mathbf{w}}\frac{1}{n}\sum_{i=1}^{n}[\psi(\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))-y_{i}\mathbf{w}^{\top}\phi(\mathbf{x}_{i})]+\lambda\|\mathbf{w}\|_{2}^{2}.$
This linearized formulation retains the nonlinear representational capacity of the kernel model while allowing efficient estimation via standard GLM solvers. The conditional mean of the response can be written as $\mathbb{E}(y|\mathbf{x})=\psi^{\prime}(\widehat{\mathbf{w}}^{\top}\phi(\mathbf{x}))$. In practice, we set $\gamma=1/p$, where $p$ is the input dimension to adapt the kernel smooth to feature scale. For the Bernoulli case, we fit a logistic regression with an $L_{2}$ penalty (inverse regularization strength $C=10$) and the lbfgs solver with a maximum of 300 iterations. For the Poisson case, we employ the same Nystrom mapping followed by a Poisson regression with $L_{2}$ regularization parameter $\alpha=10^{-3}$ and 300 iterations.
Table S.6: Simulation Results with Kernel Regression.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
@@ A.3 Additional Experiments on Kernel Regressions
$n=400,r=n^{0.8}$ -0.29(0.35) 0.37(0.27) -0.18(0.25) 0.23(0.19) 0.32 0.39 0.33 84.7% 0.83(0.46)
$r=n^{0.85}$ -0.22(0.37) 0.34(0.26) -0.13(0.27) 0.23(0.20) 0.35 0.41 0.36 89.5% 0.98(0.58)
$r=n^{0.9}$ -0.17(0.38) 0.33(0.25) -0.09(0.30) 0.23(0.21) 0.36 0.43 0.38 91.6% 1.12(0.72)
$n=700,r=n^{0.8}$ -0.20(0.28) 0.28(0.21) -0.13(0.21) 0.19(0.16) 0.26 0.34 0.27 86.7% 0.71(0.37)
$r=n^{0.85}$ -0.14(0.30) 0.26(0.20) -0.08(0.23) 0.18(0.16) 0.28 0.35 0.28 90.6% 0.81(0.45)
$r=n^{0.9}$ -0.10(0.30) 0.26(0.19) -0.05(0.24) 0.18(0.16) 0.29 0.36 0.29 91.9% 0.87(0.50)
As shown in Table S.6, the kernel based models under the ESM framework also achieve empirical coverage probabilities close to the nominal 95% level. However, compared with the random forest and deep neural network (DNN) results, the kernel approach exhibits a higher $\mathrm{MAE}_{f}$ and wider average interval lengths, suggesting that the kernel regression is less effective in capturing complex nonlinear structures than the DNN.
A.4 Additional Experiments on Dimensionality and Nonlinearity
We investigate how data dimensionality and nonlinearity influence the learning behavior of neural networks. Model performance depends not only on network capacity, but also on data quality, sample size, signal-to-noise ratio, and the complexity of the underlying regression function. We begin by assessing the effect of data dimensionality, as the inclusion of irrelevant covariates can increase noise and obscure the true signal. To quantify this effect, we compare three scenarios: low ($p=7$), moderate ($p=10$), and high ($p=20$) dimensional settings. The corresponding results are summarized in Tables S.7 and S.8.
Table S.7: Simulation results with neural networks under $p=20$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL